Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Sun Nov 16 19:27:13 MET 1997) Failure
No observations due to failure of one of the controllers of the elevation axis of the telescope.

Medicina (mauro at Mon Nov 17 06:19:44 MET 1997) Success
Total success, any scan registred with clear ski.

Westerbork (Tony Foley at Mon Nov 17 12:34:51 MET 1997) Minor failures
No phasecal, otherwise OK

Jodrell (Alastair Gunn at Mon Nov 17 17:31:16 MET 1997) Success
No known problems.

Noto (C.Trigilio at Tue Nov 18 13:41:57 MET 1997) Success
No problem reported.

Onsala (A. Polatidis at Tue Nov 18 13:43:06 MET 1997) Success
No known problems. Signal on VC01 and VC03 (LCP)

Effelsberg (dgraham at Thu Nov 20 16:42:35 MET 1997) Success

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:03:08 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name v085m.antab Read the comments in the file to get further information.